ROTATIONAL AND VIBRATIONAL ANALYSIS OF SOME LOW-LYING BENDING POLYADS IN THE Ã 1A_{u} STATE OF ACETYLENE, C_{2}H_{2}

\underline{ADAM H. STEEVES}, HANS A. BECHTEL, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139; NAMI YAMAKITA, Department of Chemical and Biological Sciences, Japan Women’s University, Mejirodai, Bunkyo-ku, Tokyo 112-8681, Japan; AN-THONY J. MERER, Institute of Atomic and Molecular Sciences, Academia Sinica, P.O. Box 23-166, Taipei, Taiwan 10617, Department of Chemistry, University of British Columbia, 6174 University Boulevard, Vancouver, B.C., Canada V6T 1Z3; and ROBERT W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139.

A number of new low-lying vibrational levels of the Ã ^{1}A_{u} state of acetylene have been identified, following infraredultraviolet double resonance experiments via the ν_{3} (Σ_{u}^{+}) and ν_{3}+ν_{4} (Π_{u}) vibrational levels of the ground state, and high sensitivity one photon laser-induced fluorescence experiments with jet-cooled samples. These new levels involve the two lowest frequency vibrations, the torsion (ν_{4}) and the in-plane bend (ν_{6}), which are nearly degenerate and have been shown to be strongly coupled by a- and b-axis Coriolis interactions^{a}. The most prominent bands in spectra recorded from the ground vibrational state or via ℓ′′ = 0 vibrational intermediates go to K′_{a} = 1 levels of the upper state (following the K′_{a} − ℓ′′ = ±1 selection rule for the transition), however data from the K′_{a} ≠ 0 levels are affected by severe a-axis Coriolis coupling, which complicates vibrational assignment. Spectra recorded from Π_{u}-symmetry vibrational intermediate states access the K′_{a} = 0 levels that reveal the purely vibrational interactions. The combinations of ν_{3} and ν_{6} are highly anharmonic. Since theoretical calculations^{b} indicate that the shape of the molecule at the cis − trans isomerization barrier will be a half-linear structure obtained by simultaneous excitation of ν_{3} and ν_{6}, this is not unexpected. The effective ν_{6} interval in the highest assigned combination (3^{5}6^{1}) is found to have dropped to 60% of the fundamental frequency, indicating that it must lie close to the barrier.

aA.L. Utz, J.D. Tobiason, E. Carrasquillo, M.L.J. Sanders and F.F. Crim, J. Chem. Phys. 95, 2742 (1993).

bJ.F. Stanton, C.M. Huang and P.G. Szalay, J. Chem. Phys. 101, 356 (1994).